# CrossMatch: Enhance Semi-Supervised Medical Image Segmentation with Perturbation Strategies and Knowledge Distillation

### v1

@@ 4.4 Ablation Study
Table 4: Ablation study of $\eta$ at all metrics on the LA dataset. 10% of labeled data are used for training in this ablation study.
$\eta$ Dice(%)↑ Jaccard(%)↑ 95HD(voxel)↓ ASD(voxel)↓
 0.1 90.88% 83.35% 7.80 1.93
 0.15 90.55% 82.92% 7.93 2.51
 0.2 91.03% 83.62% 6.44 1.99
 0.25 91.28% 84.01% 5.77 1.62
 0.3 91.33% 84.11% 5.29 1.53
 0.35 90.78% 83.21% 6.91 1.88
 0.4 87.25% 77.80% 9.80 2.37
 0.45 90.53% 82.78% 6.54 1.83
 0.5 87.78% 79.22% 6.76 1.90
@@
Table 4 displays the results of ablation experiments on the parameter of $\eta$ in the setting of 10% of the data with labels on LA dataset. The results indicate that an $\eta$ value of 0.3 yields the best performance, surpassing other values across all metrics. Consequently, we have set $\eta$ at 0.3 for all experiments in this study.
@@
Table 5: Ablation study of the performance gap between strong and weak decoders. 10% of labeled data are used for training in this ablation study.
Bottom Top Perf. Gap Dice(%)↑ Jaccard(%)↑ 95HD(voxel)↓ ASD(voxel)↓
@@
Table 5 presents the results of ablation studies on the performance gap between decoders using 10% of the data with labels on LA dataset. It is observed that the optimal model performance is achieved when the performance gap is 0.5. Notably, we also explore the scenario where the decoders are completely consistent, that is, when the performance gap is zero. In this case, the strong perturbation encoder and the weak perturbation encoder apply identical perturbations, and the model degenerates into a UniMatch with an additional $\mathcal{L}_{dkd}$.
Figure 5: Ablation study on the efficacy of various feature perturbation strategies in our method.
Fig. 5 shows the results of ablation studies on the type of Dropout used under different training sample ratios on LA dataset. We investigate three different types of Dropout available in PyTorch: standard Dropout3D, AlphaDropout, and FeatureAlphaDropout [39]. It is evident that using the standard Dropout3D as our feature perturbation strategy results in the best performance across all three data splits, followed by AlphaDropout and FeatureAlphaDropout. This may be due to the latter two inducing stronger feature perturbations, resulting in an increased Performance Gap between decoders, which is detrimental to the correct transfer of knowledge in knowledge distillation.
@@ 5 Conclusion
We have re-evaluated the role of Self-Knowledge in semi-supervised medical image segmentation and cleverly integrated feature perturbation, consistency regularization and Knowledge Distillation to propose a Self-Training segmentation method named CrossMatch. We rethink the role of perturbations in semi-supervised tasks and suggest using multiple equivalent encoders and decoders to play roles at different learning stages to expand the traditional teacher-student model, aiming to reduce the capability gap between different roles. Specifically, we derive two encoders from image-level perturbations and three decoders from feature-level perturbations, designating the unperturbed feed-forward flow as the teacher, to perform knowledge distillation on the four groups of outcomes produced by the aforementioned encoder and decoder combinations. Additionally, we utilize the properties of Mini Batches to optimize the performance of our method and provide a quantitative iteration time comparison table. Our CrossMatch demonstrate robust performance on two benchmark datasets (LA and ACDC), showing significant improvements over SOTA methods. Extensive ablation studies further validate the assumptions and design of our method.

### v2

@@ 4.5.4 Effects of Hyperparameter $\eta$
Table 11: The effects of $\eta$ at all metrics on the LA dataset. 10% of labeled data are used for training.
$\eta$ Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
 0.10 90.88 83.35 7.80 1.93
 0.15 90.55 82.92 7.93 2.51
 0.20 91.03 83.62 6.44 1.99
 0.25 91.28 84.01 5.77 1.62
 0.30 91.33 84.11 5.29 1.53
 0.35 90.78 83.21 6.91 1.88
 0.40 87.25 77.80 9.80 2.37
 0.45 90.53 82.78 6.54 1.83
 0.50 87.78 79.22 6.76 1.90
@@
Tab. 11 displays the results of experiments on the parameter of $\eta$ in the setting of 10% of the data with labels on the LA dataset. The results indicate that the $\eta$ value of 0.3 yields the best performance, surpassing other values across all metrics. Consequently, we have set $\eta$ at 0.3 for all experiments in this study.
4.5.5 Effects of Hyperparameter $\tau$
Figure 6: The effect of hyperparameter $\tau$ on performance on the LA dataset.
As shown in Fig. 6, the performance of our method on the LA dataset fluctuates with changes in the hyperparameter $\tau$. When the $\tau$ value is low, the model performs poorly, likely because it can’t effectively distinguish between correct and incorrect knowledge. As the $\tau$ value increases, the model’s performance improves, indicating that a higher $\tau$ can mitigate the transfer of incorrect knowledge. However, as $\tau$ exceeds a certain threshold, performance starts to decline, suggesting that an excessively high $\tau$ prevents the model from effectively learning from the teacher model’s knowledge. Hence we fix $\tau=0.85$ in CrossMatch.
4.5.6 Impact of Decoder Performance Gap
@@
Table 12: The performance gap between strong and weak decoders on the LA dataset. 10% of labeled data are used for training.
Bottom Top Perf. Gap Dice(%$\uparrow$ Jaccard(%$\uparrow$ 95HD(voxel)$\downarrow$ ASD(voxel)$\downarrow$
@@
Tab. 12 presents the results of the performance gap between decoders using 10% of the data with labels on the LA dataset. It is observed that the optimal model performance is achieved when the performance gap is 0.5. Notably, we also explore the scenario where the decoders are completely consistent, that is when the performance gap is zero. In this case, the strong perturbation decoder and the weak perturbation decoder apply identical perturbations, and the model degenerates into a UniMatch with an additional $\mathcal{L}_{dkd}$.
Figure 7: The efficacy of various feature perturbation strategies in our method.
4.5.7 Wrongly Segmented Examples Analysis
Figure 8: Some wrongly segmented examples using 10% labeled data on the Pancreas-CT dataset and LA dataset, respectively.
In this subsection, we examine two instances where our segmentation model demonstrates suboptimal performance, as reflected by lower Dice. These instances involve the segmentation of the Pancreas-CT dataset and the LA dataset. Through a detailed analysis of these cases, we aim to uncover the underlying causes of the segmentation errors and propose viable solutions to improve the accuracy of our model.
In the first Pancreas-CT case, the model’s Dice is 80.98%. The main challenge is insufficient segmentation of the pancreatic tail. The similarity in intensity between the pancreatic tail and adjacent tissues makes it hard to distinguish, and the small and variable structure of the pancreatic tail causes the model’s performance to be unstable in that area. In the second LA case, the model’s Dice is 87.11%. The main issue is inaccurate boundary definition caused by partial volume effects near the atrial appendages. Partial volume effects are a common problem in MRI imaging, where the signal from multiple tissue types within a single voxel gets averaged, blurring the boundaries. The anatomical complexity of the left atrium and its appendages also increases the difficulty of segmentation. Additionally, the feature-level and image-level perturbations used in our pseudo-label learning may have further interfered with the model’s ability to learn about the blurred boundaries.
To address these issues, we propose two solutions. First, advanced image processing techniques, such as multi-scale analysis or deep learning architectures that capture fine details can improve feature extraction capabilities and segmentation accuracy. However, for a fair comparison, we did not make any changes to the original V-Net model in this paper. Second, integrating advanced modelling techniques like shape constraint models or active contour models can enhance the model’s accuracy in segmenting complex structures. We are committed to continuing to optimize our method to ensure its effectiveness and reliability in clinical applications in future work.
4.5.8 Advantages and Limitations
In this study, our proposed CrossMatch shows significant performance improvement in various medical image segmentation tasks. The key advantage lies in the introduction of self-knowledge distillation and multiple perturbation strategies, which make full use of limited labeled data and a large amount of unlabeled data, greatly enhancing the model’s robustness and accuracy. Experimental results demonstrate that CrossMatch outperforms existing state-of-the-art methods on multiple benchmark datasets, particularly excelling in edge accuracy and generalization ability.
However, CrossMatch still has its limitations. Firstly, although the implementation is efficient, the computational resources required for processing large-scale 3D medical images remain high, such as the computation of multiple forward passes. Secondly, its performance still falls short in some complex structure segmentation tasks, especially in areas with high variability and fuzzy boundaries. Future research could consider incorporating more advanced image processing techniques and modelling methods to further improve the model’s accuracy and robustness.
@@ 5 Conclusion
We have re-evaluated the role of Self-Knowledge in semi-supervised medical image segmentation and cleverly integrated feature perturbation, consistency regularization and Knowledge Distillation to propose a Self-Training segmentation method named CrossMatch. We rethink the role of perturbations in semi-supervised tasks and suggest using multiple equivalent encoders and decoders to play roles at different learning stages to expand the traditional teacher-student model, aiming to reduce the capability gap between different roles. Specifically, we derive two encoders from image-level perturbations and three decoders from feature-level perturbations, designating the unperturbed feed-forward flow as the teacher to perform knowledge distillation on the four groups of outcomes produced by the aforementioned encoder and decoder combinations. Additionally, we utilize the properties of Mini Batches to optimize the performance of our method and provide a quantitative iteration time comparison table. Our CrossMatch demonstrates robust performance on four benchmark datasets (LA, ACDC, Pancreas-CT and ISIC-2018), showing significant improvements over SOTA methods. Extensive ablation studies further validate the assumptions and design of our method.